OpenStudy (anonymous):

factor completely x^5-x^3+8x^2-8

OpenStudy (anonymous):

=x^3(x^2-1)+8(x^2-1) =(x^2-1)(x^3+8) =(x-1)(x+1)(x+2)(x^2-2x+4)

OpenStudy (anonymous):

hey hw did u expand (x^3+8)?

OpenStudy (anonymous):

that's where my problem lies

OpenStudy (anonymous):

(a^3+b^3)=(a+b)(a^2-ab+b^2)

OpenStudy (anonymous):

ok thanks